OpenStudy (anonymous):

Simplify: 5(x - y) - (x - y) Question 2 options: x-y x+y 4x-4y 4x+4y

OpenStudy (anonymous):

@Brookie105

OpenStudy (anonymous):

I am not sure on this one

OpenStudy (anonymous):

can you maybe Tag someone who could help with this one? because i have no idea.. :/

OpenStudy (anonymous):

I would go with A, but I am not so sure.

OpenStudy (anonymous):

ok ill go with a

OpenStudy (anonymous):

i suppose.. if thats right

OpenStudy (anonymous):

I am not sure..

OpenStudy (freckles):

what is 5-1?

OpenStudy (anonymous):

What>?

OpenStudy (freckles):

if you have 5 apples and you take away 1 apple how many apples do you have?

OpenStudy (freckles):

are you able to take 1 away from 5? are do you need help with that?

OpenStudy (freckles):

5 take away 1 is just 4.... \[5-1=4 \\ \text{ so } 5(x-y)-1(x-y)=4(x-y) \\ \text{ now just use distributive property }\]

OpenStudy (anonymous):

um..

OpenStudy (anonymous):

can u tell me the answer and then explain.. or give me a exsample thats similar.. thats how i learn :D sorry if i seem like im being needy

OpenStudy (freckles):

distributive property goes like this: a(b+c)=a*b+a*c example 5(m+n)=5*m+5*n or can be written as 5m+5n

OpenStudy (anonymous):

i think its A x-y whatcha think?

OpenStudy (freckles):

did you actually try to use distributive property 4 doesn't equal 1... 5-1 is 4 not 1 5(x-y)-1(x-y)=4(x-y) just apply the distributive law to 4(x-y)

OpenStudy (anonymous):

um.. sorry i had lost Wifi.. one sec

OpenStudy (freckles):

multiply all the terms inside the ( ) by 4... just like I did in the example

OpenStudy (anonymous):

its either c or D i think more of C

OpenStudy (freckles):

4(x+y)=4x+4y 4(x-y)=4x-4y we had 4(x-y) so distributing 4 to the terms inside that ( ) gives 4x-4y not 4x+4y